OpenStudy (anonymous):

Find two consecutive odd integers whose product is 1443.

OpenStudy (anonymous):

guess and check

OpenStudy (anonymous):

Um, my question is how you do that.

OpenStudy (anonymous):

\(35\times 37=1295\) too small try \(37\times 39\)

OpenStudy (anonymous):

consecutive odd integers.

OpenStudy (anonymous):

Let the first odd integer be x So, the other will be : x + 2 So, x (x + 2) = 1443 \[x^2 + 2x - 1443 = 0\] Factorise it and get the answer...

OpenStudy (anonymous):

you can also do it the sissy way \[x(x+2)=1443\] which becomes \[x^2+2x-1443=0\] but you solve this by finding two integers whose product is 1443 and whose difference is 2 so it is the same problem exaclty

OpenStudy (anonymous):

@waterineyes factoring this is exactly answering the question that was asked. find two integers whose difference is 2 and whose product is 1443 making an equation out of it doesn't actually change the question in any way you have to answer the original question to answer the equation

OpenStudy (anonymous):

How do you factor this? The number's so large

OpenStudy (anonymous):

If you don't want to factorize then use the quadratic formula : D = Root(4 + 5772) = Root(5776) Root of 5776 is : 76 x = (-2 + 76)/2 or x = (-2 - 76)/2 Now can you solve it or not??

OpenStudy (anonymous):

(x+1)(x+3)=1443 SOLVE IT

OpenStudy (anonymous):

(x+1)+(x+3)= 1443